OpenStudy (anonymous):

square root of 2 sin (x-π/4)+1=0 for 0≤x≤3π

OpenStudy (xapproachesinfinity):

\(\huge \rm 2sin(x-\pi/4)+1=0\\ \huge sin(x-\pi/4)=-1/\sqrt{2}=-\sqrt{2}/2\)

OpenStudy (xapproachesinfinity):

try to solve it now

OpenStudy (xapproachesinfinity):

what angles btw 0 and 3pi give -root2/2? that what you need to find first

OpenStudy (anonymous):

135 degrees =3pi/4

OpenStudy (anonymous):

225 degrees =5pi/4

OpenStudy (anonymous):

thanks i know what to do from here

OpenStudy (xapproachesinfinity):

yes correct! good luck

OpenStudy (xapproachesinfinity):

you are welcome